Given the points A(2,5) and B(0,-3) find:
the distance between A and B
the slope of the line through A and B
The slope of a line perpendicular to the line through A and B
Please help!

Answer by venugopalramana(3286)

Given the points A(2,5)...(X1,Y1) and B(0,-3)...(X2,Y2) find:
the distance between A and B=SQRT{(X2-X1)^2+(Y2-Y1)^2}
=SQRT{(2-0)^2+(5+3)^2}=SQRT(4+64)=SQRT(68)=2*SQRT(17)

the slope of the line through A and B=(Y2-Y1)/(X2-X1)
=(-3-5)/(0-2)=4
The slope of a line perpendicular to the line through A and B=-1/4
Please help!